The larger of two positive integers is five more than twice the smaller integer. The product of the integers is 52. What are the two integers?
Answer by Greenfinch(383)

If the small integer is X, the larger is 2X + 5
So X(2X + 5) = 52
Or 2x^2 + 5X - 52 = 0
(2X + 13)(X - 4)=0
So X = -6.5 or 4
6.5 is negative and not an integer, so can be disregarded
So X = 4, larger number is 13